OpenStudy (anonymous):

How to solve: 8/26 = x/91 I know I could just type it into a calculator or Wolfram, but I need to know how to solve it because it's an essay question. My question says: What is the solution for 8/26 = x/91? Describe your steps. Do I just cross multiply since there is no x on the left hand side?

OpenStudy (ryan123345):

yes cross multiply :)

OpenStudy (anonymous):

\[\frac{ 8 }{ 26 }= \frac{ x }{ 91 }\] \[x = \frac{ 91*8 }{ 26 }\]

OpenStudy (ryan123345):

cross multiply to get 26x = 728 then u divide 26 on both sides :)

OpenStudy (anonymous):

Oh okay, gotcha. Thanks!

OpenStudy (ryan123345):

welcome :)